# Negative-Weight Single-Source Shortest Paths in Near-linear Time

### v5

@@ 1 Introduction
For the second question, note that currently all state-of-the-art results for negative-weight SSSP are based on min-cost flow algorithms, and hence rely on sophisticated continuous optimization methods and a number of complex dynamic algebraic and graph algorithms (e.g. [SW19, NSW17, ${}^{+}$20, ${}^{+}$20, NS17, Wul17]). It would be useful to develop simple efficient algorithms that are specifically tailored to negative-weight SSSP, and thus circumvent the complexity currently inherent in flow algorithms; the best known bound of this kind is still the classic $O(m\sqrt{n}\log(W))$ from over three decades ago [GT89, Gol95]. A related question is whether it is possible to achieve efficient algorithms for the problem using combinatorial tools, or whether there are fundamental barriers that make continuous optimization necessary.
@@ 1.1 Our Result
In this paper, we prioritize simplicity and modularity, and not optimizing the logarithmic factors. The follow-up work by Bringmann, Cassis, and Fischer [BCF23] greatly optimizes our framework to reduce the running time to $O(m\log^{2}(n)$ $\log(nW)\log\log n).$ There is also follow-up work by Ashvinkumar et al. showing how the framework can be applied to the parallel and distributed models of computation [${}^{+}$23].
@@ 1.2 Techniques
The decomposition above is similar to other low-diameter decompositions used in both undirected and directed graphs, though the precise guarantees vary a lot between papers [Awe85, AGLP89, AP92, ABCP92, LS93, Bar96, ${}^{+}$14, MPX13, ${}^{+}$18, FG19, CZ20, BPW20, FGdV21]. The closest similarity is to the algorithm Partition of Bernstein, Probst-Gutenberg, and Wulff-Nilsen [BPW20]. The main difference is that the algorithm of [BPW20] needed to work in a dynamic setting, and as a result their algorithm is too slow for our purposes. Our decomposition algorithm follows the general framework of [BPW20], but with several key differences to ensure faster running time; our algorithm is also simpler, since it only applies to the static setting. For the reader’s convenience, we present the entire algorithm from scratch in Section 6.
@@ 2.1 Price Functions and Equivalence
Our algorithm heavily relies on price functions, originally introduced by Johnson [Joh77]
@@ 3.2 The Interface of the Two Main Algorithms
1 // Input/Output: See Theorem 3.5 if $\Delta\leq 2$ then
@@ 2       Let $\phi_{2}=0$ and jump to Phase 3 (Algorithm 1)
3Let $d=\Delta/2$. Let $G^{B}_{\geq 0}:=(V,E,w^{B}_{\geq 0})$ where $w^{B}_{\geq 0}(e):=\max\{0,w^{B}(e)\}$ for all $e\in E$ // Phase 0: Decompose $V$ to SCCs $V_{1},V_{2},\ldots$ with weak diameter $dB$ in $G$ $E^{rem}\leftarrow\operatorname{LowDiamDecomposition}(G^{B}_{\geq 0},dB)$ (Lemma 1.2) Compute Strongly Connected Components (SCCs) of $G^{B}\setminus E^{rem}$, denoted by $V_{1},V_{2},\ldots$
@@
5 // (Lemma 4.4) If $\eta(G^{B})\leq\Delta$, then for every $v\in V_{i}$, $E[P_{G^{B}}(v)\cap E^{rem}]=O(\log^{2}n)$ // Phase 1: Make edges inside the SCCs $G^{B}[V_{i}]$ non-negative Let $H=\bigcup_{i}G[V_{i}]$, i.e. $H$ only contains edges inside the SCCs.
// (Lemma 4.5) If $G$ has no negative-weight cycle, then $\eta(H^{B})\leq d=\Delta/2$. $\phi_{1}\leftarrow\operatorname{ScaleDown}(H,\Delta/2,B)$ // (Corollary 4.6) $w_{H^{B}_{\phi_{1}}}(e)\geq 0$ for all $e\in H$ // Phase 2: Make all edges in $G^{B}\setminus E^{rem}$ non-negative $\psi\leftarrow\operatorname{FixDAGEdges}(G^{B}_{\phi_{1}}\setminus E^{rem},\{V_{1},V_{2},\ldots\})$ (Lemma 3.2) $\phi_{2}\leftarrow\phi_{1}+\psi$ // (Lemma 4.7) All edges in $(G^{B}\setminus E^{rem})_{\phi_{2}}$ are non-negative // Phase 3: Make all edges in $G^{B}$ non-negative $\psi^{\prime}\leftarrow\operatorname{ElimNeg}((G^{B}_{s})_{\phi_{2}},s)$ (Lemma 3.3) // (Theorem 4.2) expected time $O(m\log^{3}m)$. (To define $(G^{B}_{s})_{\phi_{2}}$ here, we define $\phi_{2}(s)=0$.) $\phi_{3}=\phi_{2}+\psi^{\prime}$ // (Theorem 4.1) All edges in $G^{B}_{\phi_{3}}$ are non-negative. return $\phi_{3}$ ;
@@ 4.2 Full Analysis
Note that $u$ and $v$ are in the same SCC $V_{i}$;131313in fact all vertices in $P$ are in the same SCC $V_{i}$, because we define $H=\bigcup_{i}G[V_{i}]$. thus, by Lemma 4.3:
@@
1 $\bar{w}(e)\leftarrow w_{in}(e)\cdot 2n$ for all $e\in E$, $\bar{G}\leftarrow(V,E,\bar{w})$, $B\leftarrow 2n$. // scale up edge weights Round $B$ up to nearest power of 2 // still have $\bar{w}(e)\geq-B$ for all $e\in E$ $\phi_{0}(v)=0$ for all $v\in V$ // identity price function for $i=1$ to $t:=\log_{2}(B)$ do
2       $\psi_{i}\leftarrow\operatorname{ScaleDown}(\bar{G}_{\phi_{i-1}},\Delta:=n,B/2^{i})$ $\phi_{i}\leftarrow\phi_{i-1}+\psi_{i}$ // (Claim 5.1) $w_{\phi_{i}}(e)\geq-B/2^{i}$ for all $e\in E$
$G^{*}\leftarrow(V,E,w^{*})$ where $w^{*}(e)\leftarrow\bar{w}_{\phi_{t}}(e)+1$ for all $e\in E$. // Observe: $G^{*}$ in above line has non-negative weights Compute a shortest path tree $T$ from $s$ using $\operatorname{Dijkstra}(G^{*},s)$ (Lemma 3.1) // (5.3) Will Show: any shortest path in $G^{*}$ is also shortest in $G$ return shortest path tree $T$
@@ 6 Algorithm for Low-Diameter Decomposition
Given a directed graph $G=(V,E)$, a vertex $v\in V$, and a distance-parameter $R$, we define $\operatorname{Ball_{G}^{out}}(v,R)=\{u\in V\mid\operatorname{dist}(v,u)\leq R\}$. We define ${\sf boundary}(\operatorname{Ball_{G}^{out}}(v,R))=\{(x,y)\in E\mid x\in\operatorname{Ball_{G}^{out}}(v,R)\land y\notin\operatorname{Ball_{G}^{out}}(v,R)\}$. Similarly, we define $\operatorname{Ball_{G}^{in}}(v,R)=\{u\in V\mid\operatorname{dist}(u,v)\leq R\}$ and we define ${\sf boundary}(\operatorname{Ball_{G}^{in}}(v,R))=\{(x,y)\in E\mid y\in\operatorname{Ball_{G}^{in}}(v,R)\land x\notin\operatorname{Ball_{G}^{in}}(v,R)\}$. We often use $\operatorname{Ball_{G}^{*}}$ to denote that a ball can be either an out-ball or an in-ball, i.e., *$*\in\{in,out\}$.
@@
1 Let $n$ be the global variable in Remark 6.3 $G_{0}\leftarrow G,E^{rem}\leftarrow\emptyset$ // Phase 1: mark vertices as light or heavy $k\leftarrow c\ln(n)$ for large enough constant $c$ $S\leftarrow\{s_{1},\ldots,s_{k}\}$, where each $s_{i}$ is a random node in $V$ // possible: $s_{i}=s_{j}$ for $i\neq j$ For each $s_{i}\in S$ compute $\operatorname{Ball_{G}^{in}}(s_{i},D/4)$ and $\operatorname{Ball_{G}^{out}}(s_{i},D/4)$ For each $v\in V$ compute $\operatorname{Ball_{G}^{in}}(v,D/4)\bigcap S$ and $\operatorname{Ball_{G}^{out}}(v,D/4)\bigcap S$ using Line 3 foreach $v\in V$ do
2       If $|\operatorname{Ball_{G}^{in}}(v,D/4)\bigcap S|\leq.6k$, mark $v$ in-light // whp $|\operatorname{Ball_{G}^{in}}(v,D/4)|\leq.7|V(G)|$ Else if $|\operatorname{Ball_{G}^{out}}(v,D/4)\bigcap S|\leq.6k$, mark $v$ out-light // whp $|\operatorname{Ball_{G}^{out}}(v,D/4)|\leq.7|V(G)|$ Else mark $v$ heavy // w.h.p $\operatorname{Ball_{G}^{in}}(v,D/4)>.5|V(G)|$ and $\operatorname{Ball_{G}^{out}}(v,D/4)>.5|V(G)|$
3// Phase 2: carve out balls until no light vertices remain while $G$ contains a node v marked ***-light for *$*\in\{in,out\}$ do
4       Sample $R_{v}\sim\operatorname{Geo}(p)$ for $p=\min\{1,80\log_{2}(n)/D\}$. Compute $\operatorname{Ball_{G}^{*}}(v,R_{v})$. $E^{boundary}\leftarrow{\sf boundary}(\operatorname{Ball_{G}^{*}}(v,R_{v}))$ // add boundary edges of ball to $E^{rem}$. If $R_{v}>D/4$ or $|\operatorname{Ball_{G}^{*}}(v,R_{v})|>.7|V(G)|$ then return $E^{rem}\leftarrow E(G)$ and terminate // Pr[terminate] $\leq 1/n^{20}$ $E^{recurse}\leftarrow\operatorname{LDD}(G[\operatorname{Ball_{G}^{*}}(v,R_{v})],D)$ // recurse on ball $E^{rem}\leftarrow E^{rem}\bigcup E^{boundary}\bigcup E^{recurse}$. $G\leftarrow G\setminus\operatorname{Ball_{G}^{*}}(v,R_{v})$ // remove ball from $G$
// Clean Up: check that remaining vertices have small weak diameter in initial input graph $G_{0}$. Select an arbitrary vertex $v$ in $G$. If $\operatorname{Ball}_{G_{0}}^{in}(v,D/2)\nsupseteq V(G)$ or $\operatorname{Ball}_{G_{0}}^{out}(v,D/2)\nsupseteq V(G)$ then return $E^{rem}\leftarrow E(G)$ and terminate // Pr[terminate] $\leq 1/n^{20}$. If this does not terminate, then all remaining vertices in $V(G)$ have weak diameter $\leq D$ Return $E^{rem}$
@@
The pseudocode for our low-diameter decomposition algorithm can be found in Algorithm 3. Roughly, in Phase 1 each vertex is marked as either in-light, out-light, or heavy. In Phase 2 we repeatedly “remove” balls centered at either in-light or out-light vertices: Let $v$ be any in-light vertex (the process is similar for out-light vertices). Consider a ball $\operatorname{Ball_{G}^{in}}(v,R_{v})$ with radius $R_{v}$ selected randomly from the geometric distribution $\operatorname{Geo}(p)$, with $p=\min\{1,80\log(n)/D\}$ (Algorithms 3 and 3) [^16] Throughout this section, $\log(n)$ always denotes $\log_{2}(n)$. We add edges pointing into this ball (i.e. edges in ${\sf boundary}(\operatorname{Ball_{G}^{in}}(v,R_{v}))$) to $E^{boundary}$, which will be later added to $E^{rem}.$ We recurse the algorithm on this ball (Algorithm 3) which may add more edges to $E^{rem}$ (via $E^{recurse}$). Finally, we remove this ball from the graph and repeat the process. Note that the algorithm may also terminate prematurely with $E^{rem}=E(G)$ in Algorithms 3 and 3. We will show that this happens with very low probability (Section 6.3).
@@ 6.2 Running Time Analysis
Every time the algorithm executes the While loop in Line 3 it first samples $R_{v}\sim\operatorname{Geo}(p)$ from the geometric distribution; this can be done in $O(\log(n))$ time using e.g. [BF13]. [^17] By Theorem 5 in [BF13] with $p=\log(n)/D$ and word size $w=\Omega(\log\log D)$, the expected time is $O(\log(\min\{D/\log n,n\})/\log w)=O(\log n).$ Note that there might be other simpler methods since we do not need an exact algorithm. It then computes some ball $\operatorname{Ball_{G}^{*}}(v,R_{v})$ (Algorithm 3). All vertices in this ball are then removed from $G$ (Line 3), so at this level of recursion each vertex participates in at most one ball. Each ball is computed via Dijkstra’s algorithm, which requires $O(\log(n))$ time per vertex explored and $O(1)$ time per edge explored.181818 A standard Dijkstra’s implementation requires $O(|V(G)|)$ time to initialize the priority queue and vertices’ labels (e.g. [CLRS09]). This can be easily modified to avoid the initialization time (e.g. [TZ05]). One way to do this is to initialize the priority queue and labels once in the beginning of Phase 2. After we use them to compute each ball $\operatorname{Ball_{G}^{*}}(v,R_{v})$ on Algorithm 3, we reinitialize them at the cost of the number of explored vertices. So, the total time to execute the while loop is $O(|V(G)|\log(n)+|E(G)|)$.

### v6

@@ 1 Introduction
For the second question, note that currently all state-of-the-art results for negative-weight SSSP are based on min-cost flow algorithms, and hence rely on sophisticated continuous optimization methods and a number of complex dynamic algebraic and graph algorithms (e.g. [SW19, NSW17, CGL+20, BBP+20, NS17, Wul17]). It would be useful to develop simple efficient algorithms that are specifically tailored to negative-weight SSSP, and thus circumvent the complexity currently inherent in flow algorithms; the best known bound of this kind is still the classic $O(m\sqrt{n}\log(W))$ from over three decades ago [GT89, Gol95]. A related question is whether it is possible to achieve efficient algorithms for the problem using combinatorial tools, or whether there are fundamental barriers that make continuous optimization necessary.
@@ 1.1 Our Result
In this paper, we prioritize simplicity and modularity, and not optimizing the logarithmic factors. The follow-up work by Bringmann, Cassis, and Fischer [BCF23] greatly optimizes our framework to reduce the running time to $O(m\log^{2}(n)$ $\log(nW)\log\log n).$ There is also follow-up work by Ashvinkumar et al. showing how the framework can be applied to the parallel and distributed models of computation [ABC+23].
@@ 1.2 Techniques
The decomposition above is similar to other low-diameter decompositions used in both undirected and directed graphs, though the precise guarantees vary a lot between papers [Awe85, AGLP89, AP92, ABCP92, LS93, Bar96, BGK+14, MPX13, PRS+18, FG19, CZ20, BPW20, FGdV21]. The closest similarity is to the algorithm Partition of Bernstein, Probst-Gutenberg, and Wulff-Nilsen [BPW20]. The main difference is that the algorithm of [BPW20] needed to work in a dynamic setting, and as a result their algorithm is too slow for our purposes. Our decomposition algorithm follows the general framework of [BPW20], but with several key differences to ensure faster running time; our algorithm is also simpler, since it only applies to the static setting. For the reader’s convenience, we present the entire algorithm from scratch in Section 6.
@@ 2.1 Price Functions and Equivalence
Our algorithm heavily relies on price functions—an idea with deep roots, traceable to Dreyfus’ 1969 survey [Dre69] and Braess’ work [Bra71, Page 176], if not earlier—and which also appeared in the well-known $A^{*}$ algorithm (referred to as a heuristic function in the AI community). [^10] We thank Donald E. Knuth for the historical note regarding price functions.
@@ 3.2 The Interface of the Two Main Algorithms
1
2// Input/Output: See Theorem 3.5
3if $\Delta\leq 2$ then
@@ 4       Let $\phi_{2}=0$ and jump to Phase 3 (Algorithm 1)
5
6Let $d=\Delta/2$. Let $G^{B}_{\geq 0}:=(V,E,w^{B}_{\geq 0})$ where $w^{B}_{\geq 0}(e):=\max\{0,w^{B}(e)\}$ for all $e\in E$
7// Phase 0: Decompose $V$ to SCCs $V_{1},V_{2},\ldots$ with weak diameter $dB$ in $G$
[…]
9Compute Strongly Connected Components (SCCs) of $G^{B}\setminus E^{rem}$, denoted by $V_{1},V_{2},\ldots$
@@
11 // (Lemma 4.4) If $\eta(G^{B})\leq\Delta$, then for every $v\in V_{i}$, $E[P_{G^{B}}(v)\cap E^{rem}]=O(\log^{2}n)$
12// Phase 1: Make edges inside the SCCs $G^{B}[V_{i}]$ non-negative
13Let $H=\bigcup_{i}G[V_{i}]$, i.e. $H$ only contains edges inside the SCCs.
14 // (Lemma 4.5) If $G$ has no negative-weight cycle, then $\eta(H^{B})\leq d=\Delta/2$.
[…]
16// Phase 2: Make all edges in $G^{B}\setminus E^{rem}$ non-negative
[…]
19// Phase 3: Make all edges in $G^{B}$ non-negative
[…]
22
return $\phi_{3}$ ;
@@ 4.2 Full Analysis
Note that $u$ and $v$ are in the same SCC $V_{i}$;141414in fact all vertices in $P$ are in the same SCC $V_{i}$, because we define $H=\bigcup_{i}G[V_{i}]$. thus, by Lemma 4.3:
@@
1
[…]
3Round $B$ up to nearest power of 2 // still have $\bar{w}(e)\geq-B$ for all $e\in E$
[…]
5for $i=1$ to $t:=\log_{2}(B)$ do
6
7      $\psi_{i}\leftarrow\operatorname{ScaleDown}(\bar{G}_{\phi_{i-1}},\Delta:=n,B/2^{i})$
8      $\phi_{i}\leftarrow\phi_{i-1}+\psi_{i}$ // (Claim 5.1) $w_{\phi_{i}}(e)\geq-B/2^{i}$ for all $e\in E$
9
[…]
11// Observe: $G^{*}$ in above line has non-negative weights
12Compute a shortest path tree $T$ from $s$ using $\operatorname{Dijkstra}(G^{*},s)$ (Lemma 3.1)
13// (5.3) Will Show: any shortest path in $G^{*}$ is also shortest in $G$
14return shortest path tree $T$
@@ 6 Algorithm for Low-Diameter Decomposition
Given a directed graph $G=(V,E)$, a vertex $v\in V$, and a distance-parameter $R$, we define $\operatorname{Ball_{G}^{out}}(v,R)=\{u\in V\mid\operatorname{dist}(v,u)\leq R\}$. We define $\partial{\operatorname{Ball_{G}^{out}}(v,R)}=\{(x,y)\in E\mid x\in\operatorname{Ball_{G}^{out}}(v,R)\land y\notin\operatorname{Ball_{G}^{out}}(v,R)\}$. Similarly, we define $\operatorname{Ball_{G}^{in}}(v,R)=\{u\in V\mid\operatorname{dist}(u,v)\leq R\}$ and we define $\partial{\operatorname{Ball_{G}^{in}}(v,R)}=\{(x,y)\in E\mid y\in\operatorname{Ball_{G}^{in}}(v,R)\land x\notin\operatorname{Ball_{G}^{in}}(v,R)\}$. We often use $\operatorname{Ball_{G}^{*}}$ to denote that a ball can be either an out-ball or an in-ball, i.e., $*\in\{in,out\}$.
@@
1
2Let $n$ be the global variable in Remark 6.3
[…]
4// Phase 1: mark vertices as light or heavy
[…]
7For each $s_{i}\in S$ compute $\operatorname{Ball_{G}^{in}}(s_{i},D/4)$ and $\operatorname{Ball_{G}^{out}}(s_{i},D/4)$
8For each $v\in V$ compute $\operatorname{Ball_{G}^{in}}(v,D/4)\bigcap S$ and $\operatorname{Ball_{G}^{out}}(v,D/4)\bigcap S$ using Line 3
9foreach $v\in V$ do
10
11      If $|\operatorname{Ball_{G}^{in}}(v,D/4)\bigcap S|\leq.6k$, mark $v$ in-light // whp $|\operatorname{Ball_{G}^{in}}(v,D/4)|\leq.7|V(G)|$
12      Else if $|\operatorname{Ball_{G}^{out}}(v,D/4)\bigcap S|\leq.6k$, mark $v$ out-light // whp $|\operatorname{Ball_{G}^{out}}(v,D/4)|\leq.7|V(G)|$
13      Else mark $v$ heavy // w.h.p $\operatorname{Ball_{G}^{in}}(v,D/4)>.5|V(G)|$ and $\operatorname{Ball_{G}^{out}}(v,D/4)>.5|V(G)|$
14
15// Phase 2: carve out balls until no light vertices remain
16while $G$ contains a node v marked *-light for $*\in\{in,out\}$ do
17
18      Sample $R_{v}\sim\operatorname{Geo}(p)$ for $p=\min\{1,80\log_{2}(n)/D\}$.
19      Compute $\operatorname{Ball_{G}^{*}}(v,R_{v})$.
20      $E^{boundary}\leftarrow\partial{\operatorname{Ball_{G}^{*}}(v,R_{v})}$ // add boundary edges of ball to $E^{rem}$.
21      If $R_{v}>D/4$ or $|\operatorname{Ball_{G}^{*}}(v,R_{v})|>.7|V(G)|$ then return $E^{rem}\leftarrow E(G)$ and terminate // Pr[terminate] $\leq 1/n^{20}$
22      $E^{recurse}\leftarrow\operatorname{LDD}(G[\operatorname{Ball_{G}^{*}}(v,R_{v})],D)$ // recurse on ball
23      $E^{rem}\leftarrow E^{rem}\bigcup E^{boundary}\bigcup E^{recurse}$.
24      $G\leftarrow G\setminus\operatorname{Ball_{G}^{*}}(v,R_{v})$ // remove ball from $G$
25
26// Clean Up: check that remaining vertices have small weak diameter in initial input graph $G_{0}$.
27Select an arbitrary vertex $v$ in $G$.
28If $\operatorname{Ball}_{G_{0}}^{in}(v,D/2)\nsupseteq V(G)$ or $\operatorname{Ball}_{G_{0}}^{out}(v,D/2)\nsupseteq V(G)$ then return $E^{rem}\leftarrow E(G)$ and terminate // Pr[terminate] $\leq 1/n^{20}$. If this does not terminate, then all remaining vertices in $V(G)$ have weak diameter $\leq D$
29Return $E^{rem}$
@@ 24      $G\leftarrow G\setminus\operatorname{Ball_{G}^{*}}(v,R_{v})$ // remove ball from $G$
The pseudocode for our low-diameter decomposition algorithm can be found in Algorithm 3. Roughly, in Phase 1 each vertex is marked as either in-light, out-light, or heavy. In Phase 2 we repeatedly “remove” balls centered at either in-light or out-light vertices: Let $v$ be any in-light vertex (the process is similar for out-light vertices). Consider a ball $\operatorname{Ball_{G}^{in}}(v,R_{v})$ with radius $R_{v}$ selected randomly from the geometric distribution $\operatorname{Geo}(p)$, with $p=\min\{1,80\log(n)/D\}$ (Algorithms 3 and 3) [^17] Throughout this section, $\log(n)$ always denotes $\log_{2}(n)$. We add edges pointing into this ball (i.e. edges in $\partial{\operatorname{Ball_{G}^{in}}(v,R_{v})}$) to $E^{boundary}$, which will be later added to $E^{rem}.$ We recurse the algorithm on this ball (Algorithm 3) which may add more edges to $E^{rem}$ (via $E^{recurse}$). Finally, we remove this ball from the graph and repeat the process. Note that the algorithm may also terminate prematurely with $E^{rem}=E(G)$ in Algorithms 3 and 3. We will show that this happens with very low probability (Section 6.3).
@@ 6.2 Running Time Analysis
Every time the algorithm executes the While loop in Line 3 it first samples $R_{v}\sim\operatorname{Geo}(p)$ from the geometric distribution; this can be done in $O(\log(n))$ time using e.g. [BF13]. [^18] By Theorem 5 in [BF13] with $p=\log(n)/D$ and word size $w=\Omega(\log\log D)$, the expected time is $O(\log(\min\{D/\log n,n\})/\log w)=O(\log n).$ Note that there might be other simpler methods since we do not need an exact algorithm. It then computes some ball $\operatorname{Ball_{G}^{*}}(v,R_{v})$ (Algorithm 3). All vertices in this ball are then removed from $G$ (Line 3), so at this level of recursion each vertex participates in at most one ball. Each ball is computed via Dijkstra’s algorithm, which requires $O(\log(n))$ time per vertex explored and $O(1)$ time per edge explored.191919 A standard Dijkstra’s implementation requires $O(|V(G)|)$ time to initialize the priority queue and vertices’ labels (e.g. [CLRS09]). This can be easily modified to avoid the initialization time (e.g. [TZ05]). One way to do this is to initialize the priority queue and labels once in the beginning of Phase 2. After we use them to compute each ball $\operatorname{Ball_{G}^{*}}(v,R_{v})$ on Algorithm 3, we reinitialize them at the cost of the number of explored vertices. So, the total time to execute the while loop is $O(|V(G)|\log(n)+|E(G)|)$.